OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−7, 5) and a directrix of y = −11

OpenStudy (gnorris):

f(x) = one thirty second (x + 7)2 − 3

OpenStudy (gnorris):

(x - x)^2 + (y - -11)^2 = (x - -7)^2 + (y - 5)^2 y^2 + 22y + 121 = (x + 7)^2 + y^2 -10y + 25 32 y = (x + 7)^2 -96 y = (1/32)(x + 7)^2 - 3

OpenStudy (anonymous):

@gnorris thank you

OpenStudy (gnorris):

you're welcome